OpenStudy (anonymous):

The circle x^2 + y^2 - 4x - 2y = 4 has two x-intercepts x1 and x2 as well as two y-intercepts y1 and y2. Verify the following:

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

a) x1 + x2 = 4

OpenStudy (anonymous):

b) y1 + y2 = 2

OpenStudy (anonymous):

c) |x1x2|=|y1y2|=4

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

help me @perl

OpenStudy (anonymous):

please! ive been stuck for a long time

OpenStudy (perl):

you can find the equation of a circle (x-h)^2 + (y - k)^2 = r^2

OpenStudy (anonymous):

i did that already

OpenStudy (perl):

what do you have

OpenStudy (perl):

the x intercepts occur when y = 0

OpenStudy (anonymous):

i got (x-2)^2 + (y-1)^2 = 9

OpenStudy (anonymous):

any ideas wio?

OpenStudy (anonymous):

Well, when you have an \(x\) intercept \(x_1\), that means there is a point \((x_1,0)\).

OpenStudy (anonymous):

So one thing you can do is plug in that point, I suppose.

OpenStudy (anonymous):

into the equation of the circle?

OpenStudy (anonymous):

Into the equation they gave you.

OpenStudy (anonymous):

For example, for the point \((x_1,0)\), we have:\[ (x_1)^2 + (0)^2 - 4(x_1) - 2(0) = 4 \]

OpenStudy (triciaal):

@Andrew12840 i got (x-2)^2 + (y-1)^2 = 9 are you sure its =9 and not = 10

OpenStudy (anonymous):

ok so i got (a), is (b) the same thing?

OpenStudy (anonymous):

its not working for b)

OpenStudy (anonymous):

i get y^2-2y-4=0

OpenStudy (anonymous):

Okay, I think you're a little confused

OpenStudy (anonymous):

I'm basically thinking we solve for the intercepts.

OpenStudy (anonymous):

\[ x^2 + y^2 - 4x - 2y = 4 \]If we let \(y=0\), then: \[ x^2-4x-4=0 \]Using the quadratic formula will give use our result.

OpenStudy (anonymous):

It will give us our two intercepts

OpenStudy (anonymous):

i understand that im talking about problem b) y1+y2=2

OpenStudy (anonymous):

For that we'd have: \[ y^2-2y-4=0 \]

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

where do i go from there

OpenStudy (anonymous):

what did you get for the roots?

OpenStudy (anonymous):

1 plus or minus rad 5

OpenStudy (triciaal):

got the same

OpenStudy (anonymous):

Add and you will get \(2\).

OpenStudy (anonymous):

The \(\pm\) part cancels out

OpenStudy (anonymous):

what?

OpenStudy (triciaal):

b) verify y1 + y2 = 2 left hand side y1 = 1 + rad 5 and y2 = 1 - rad 5 y1 + y2 = 1 + rad 5 + 1 - rad 5 = 2 = right hand side

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

do you know how to do c)?

OpenStudy (triciaal):

repost it here

OpenStudy (anonymous):

c) |x1x2|=|y1y2|=4

OpenStudy (anonymous):

i have to verify it

OpenStudy (anonymous):

Since you already know all the intercepts, just plug it in.

OpenStudy (anonymous):

thanks to both of you

OpenStudy (triciaal):

you are welcome